OpenStudy (anonymous):

Help? The ratio of butter:flour:sugar in a recipe is 1:9:6. Using 8 cups of sugar, how many cups will be used in this recipe( total ) ?

OpenStudy (anonymous):

I'll fan anyone and medal who answers

Directrix (directrix):

What I want to know is if the numbers of cups of the integers have to be whole numbers?

Directrix (directrix):

@SmartiesNerd Is one of your answer options a fraction or mixed number?

OpenStudy (anonymous):

It's a ratio and yes they have to be whole numebrs

Directrix (directrix):

Did you state the problem exactly as it appears in the text? Because the problem doesn't say that the cups have to be whole numbers. So, how do you know? I need to know so I can help with the problem.

OpenStudy (anonymous):

Oh well what I wrote is exactly what the problem stated and since we're learning ratios I assumed they would be whole numbers.

Directrix (directrix):

If you use 8 cups of sugar, then to figure out how to maintain the ratios, solve 6x = 8 and get 8/6 or 4/3 as the common factor to maintain the ratios.

Directrix (directrix):

What is 4/3 * 9 = @SmartiesNerd

OpenStudy (anonymous):

I think its 12.

Directrix (directrix):

B F S 1 :9 :6 4/3 12 : 8

OpenStudy (anonymous):

Thanks for helping me btw

Directrix (directrix):

Let's see if that works.

OpenStudy (anonymous):

omg. I just noticed that it says to express your answer as a common fraction

OpenStudy (anonymous):

I'm so sorry I didn't notice that before

Directrix (directrix):

9/6 = 12/8 ? 9/12 = 1/(4/3) ? 1/(4/3) = 6/8 ? We have to check these to be sure we kept that original ratio.

Directrix (directrix):

They are good.

OpenStudy (anonymous):

Thank you thank you thank youu

Directrix (directrix):

Add these: 4/3 + 12 + 8 = ? @SmartiesNerd

OpenStudy (anonymous):

21.3( continuing )

Directrix (directrix):

Yes. Call it 21 1/3 because the answer has to be a fraction.

OpenStudy (anonymous):

oh ok

Directrix (directrix):

So, multiply 21 * 3. To that add 1 --> that sum will be the numerator of the fraction The denominator will be 3.

OpenStudy (anonymous):

Thanks alot for helping me

Directrix (directrix):

You're welcome but first, what did you get for the final answer?

Directrix (directrix):

@SmartiesNerd The answer ?

OpenStudy (anonymous):

I got 21 1/3 and when I entered it online it said I was correct but they had it in a mixed fraction form

Directrix (directrix):

That is what we were doing when you left me. >>. So, multiply 21 * 3. To that add 1 64/3

OpenStudy (anonymous):

Thanksss i'm sorry it's just i'm a slow learner

Directrix (directrix):

I don't agree. You are welcome for the help.